# Cross-Modal Causal Intervention for Medical Report Generation

### v4

@@ Abstract
Medical report generation (MRG) is essential for computer-aided diagnosis and medication guidance, which can relieve the heavy burden of radiologists by automatically generating the corresponding medical reports according to the given radiology image. However, due to the spurious correlations within image-text data induced by visual and linguistic biases, it is challenging to generate accurate reports reliably describing lesion areas. Moreover, the cross-modal confounders are usually unobservable and challenging to be eliminated explicitly. In this paper, we aim to mitigate the cross-modal data bias for MRG from a new perspective, i.e., cross-modal causal intervention, and propose a novel Visual-Linguistic Causal Intervention (VLCI) framework for MRG, which consists of a visual deconfounding module (VDM) and a linguistic deconfounding module (LDM), to implicitly mitigate the visual-linguistic confounders by causal front-door intervention. Specifically, due to the absence of a generalized semantic extractor, the VDM explores and disentangles the visual confounders from the patch-based local and global features without expensive fine-grained annotations. Simultaneously, due to the lack of knowledge encompassing the entire field of medicine, the LDM eliminates the linguistic confounders caused by salient visual features and high-frequency context without constructing a terminology database. Extensive experiments on IU-Xray and MIMIC-CXR datasets show that our VLCI significantly outperforms the state-of-the-art MRG methods. The code and models are available at https://github.com/WissingChen/VLCI.
@@
Medical report generation, causality, visual-language pre-training model.I Introduction
@@
Medical images (e.g., X-Ray, MRI) are widely used in clinical procedures [1], providing significant evidence for disease analysis and medical diagnosis [2]. Nevertheless, observing suspicious lesions and writing a coherent diagnosis report is time-consuming, even for experienced radiologists. Furthermore, inexperienced radiologists often fail to capture tiny abnormalities due to the high requirement for clinical knowledge. To relieve this issue, Medical Report Generation (MRG) has emerged and attracted growing interest in recent years [3]. MRG extracts features from medical images and generates the corresponding reports, which is similar to image captioning [4]. However, the current MRG faces three challenges that are significantly different from the image captioning task: 1) Longer sentence generation (60-100 tokens) tends to accumulate a larger bias, whereas image captions typically consist of fewer than 20 tokens [5]. 2) The necessity to capture all key regions in the medical image (i.e., abnormalities and diagnostic evidence) and a low tolerance for factually incorrect information [3]. 3) More sophisticated linguistic and visual semantic patterns require a proficient understanding of complex medical information, whereas entities in natural images are diverse and easily distinguishable [6, 2]. Therefore, these challenges impose significant limitations on modeling visual-linguistic interactions and learning informative cross-modal representations for accurate MRG [5].
@@
To tackle the aforementioned challenges, current MRG methods have made significant efforts, such as the memory-driven module for longer sentence generation [5], additional knowledge for more accurate description [7], and contrast with normal samples (i.e., images without lesion areas) for the capture of abnormalities (i.e., lesion areas within images) [8]. Actually, most of the previous MRG methods aim to capture the latent subtle differences in images (visual biases) and learn a concise set of key descriptions in the text (linguistic biases) for accurate long-sequence generation. Moreover, these methods usually focus on training computationally expensive models based on a large number of region-level annotations [3] and task-specific knowledge111These methods build the template or knowledge database laboriously, making it hard to transfer those approaches directly to other datasets [9]., rather than mitigating cross-modal biases. Nevertheless, significant visual and linguistic biases exist in numerous image-text data, as shown in Fig. 1. Therefore, lightweight models that can mitigate the cross-modal data bias are essential for MRG to accurately discover abnormalities and generate reliable reports.
Moreover, the most significant difficulty in MRG is the existence of visual and linguistic biases that lead to entangled cross-modal features, i.e., spurious correlations, causing the incorrect report in the model prediction with confounders. This means that the complexity of the medical diagnosis requires the discovery of true causal relations rather than spurious correlations, especially for medical diagnosis when considering rare cases. Actually, the majority of samples in medical reports consist of normal individuals, while there are a limited number of abnormal samples that require comprehensively considering various organs to determine pathological changes in other regions. Additionally, there exist significant visual-linguistic biases in the training sets, which makes the MRG model predict wrong reports in the testing set when relying on the spurious correlations in the training set, as shown in Fig. 1(a-b). This indicates the necessity to uncover true causal relations. Otherwise, spurious correlations may be employed, leading to an inaccurate estimation of “No Effusion” instead of “Pleural Effusion”, as shown in Fig. 1(c-d). Specifically, visual confounders can be regarded as visual features related to the cardiac region; linguistic confounders pertain to descriptions of a normal heart; and cross-modal confounders arise due to the emphasis on visual features of the cardiac region, leading to wrong judgments.
@@
To mitigate the visual-linguistic bias, causal inference [10, 11] has shown promising performance in image classification [12], image question answering [13], and image captioning [14]. However, directly applying existing causal methods to the MRG task may yield unsatisfactory results due to the unobservable confounders in the visual and linguistic domain and the complex visual-linguistic interaction in medical images and textual reports. Although back-door intervention can cut off the shortcut path, it requires approximating the observable confounders using a well-trained visual object detector or a well-constructed linguistic dictionary. Fortunately, causal front-door intervention gives a feasible way to calculate unobservable confounders and mitigate visual-linguistic spurious correlations. With causal front-door intervention, we can eliminate the spurious cross-modal correlations effectively by introducing an additional mediator [15, 16] and generate an accurate description of “Pleural Effusion”. The mediator can be assumed as the feature of “Pleural” in different findings (e.g., “Normal Heart”) to estimate the sub-distribution of “Pleural Effusion” [17]. However, the accurate and reasonable acquisition of the mediator is challenging, especially without the support of additional medical knowledge.
@@
Motivated by the effectiveness of causal inference in deconfounding the cross-modal bias, we propose a lightweight cross-modal causal intervention framework for MRG without the observable confounders assumption, named Visual-Linguistic Causal Intervention (VLCI), to mitigate the visual and linguistic data biases. Specifically, we propose Prefix Language Modeling (PLM) and degradation-aware Masked Image Modeling (MIM) for cross-modal pre-training. To mitigate the visual and linguistic biases, we propose the visual deconfounding module (VDM) and linguistic deconfounding module (LDM) based on the causal front-door intervention paradigm. In VDM, the visual mediator is constructed by local detail information (e.g., lung texture) and global contour (e.g., lung contour) from medical images, to disentangle the visual features. The linguistic confounders can be eliminated by the LDM, which estimates the change in the probability of word embedding caused by visual details and linguistic context. In summary, our main contributions are listed as follows:
To implicitly mitigate cross-modal biases, we propose visual-linguistic causal front-door intervention modules VDM and LDM based on the Structural Causal Model. The VDM aims to disentangle the region-based features from images, and the LDM aims to eliminate the spurious correlations caused by the visual-linguistic embedding.
@@
To alleviate the problem of unpaired data and capture detailed features when pre-training cross-modal data, we propose PLM and degradation-aware MIM for pre-training in various data situations (e.g., unpaired, single modality), which is efficient and easy to implement.
@@
We propose a lightweight Visual-Linguistic Causal Intervention (VLCI) framework for MRG, which introduces mediators without additional knowledge, to implicitly deconfound the visual-linguistic features by causal front-door intervention. To the best of our knowledge, we are the first to conduct cross-modal causal intervention for MRG. Experiments show that we achieve state-of-the-art performance on IU-Xray and MIMIC-CXR datasets.
@@ II-A Image Captioning
Image captioning aims to understand image information and describe it in text, which mainly adopts the encoder-decoder framework [18]. Recent work has achieved significant success in this task, generally improving model performance from three aspects: visual representation learning, linguistic comprehension and generation, and training strategy [19]. The global features extractor from CNNs leads to excessive compression of information and lacks granularity [20], thus, visual saliency can further improve performance via the spatial relation of regional features and rely on the integration of visual and semantic data to improve performance [21, 22]. As for the linguistics features, generating the captioning from coarse to fine [23] and integrating multi-modal information via multiple layers of LSTM [24] achieve a promising performance, but are limited by the training efficiency and expression ability. With the rise of transformer-based models [18], intrinsic information of each modality is learned through self-attention, while cross-attention enables the learning of multi-modal information, thereby effectively enhancing the model’s performance [25, 14]. Furthermore, the performance improvements can be achieved by the formulation of training strategies, including the learning order of the samples [26], reinforcement learning [4], and visual-linguistic pre-training [27, 28]. Compared with the image captioning approaches, the MRG has similar structures [19]. Nevertheless, image captioning usually generates a single sentence to describe the main entities, while the MRG focuses on potential subtle lesion areas in medical images and generates longer sentences from more sophisticated visual-linguistic semantics.
@@ II-B Causal Inference
As the visual-linguistic bias caused by the heterogeneity of multi-modal data, causality provides a new methodology to design robust models via the elimination of spurious correlation [10, 29, 11]. Causal inference estimates the hidden causal effects in the distribution while significantly improving the model’s generalization. It mitigates confounders through back-door, front-door intervention, or counterfactual intervention. This approach substantially advances the performance of tasks, including image classification [12], image semantic segmentation [30], visual features representation [31, 32], image captioning [17, 14], and VQA [15]. Specifically, Wang et al. [13] improved Faster R-CNN by causal back-door intervention to obtain a more robust object detection model, which improves the performance of VQA [33] and image captioning [14]. Furthermore, it can also achieve deconfounding by simulating causal interventions based on human priors in modeling [34].
However, the confounders are usually unobservable and elusive, thus front-door intervention and counterfactual intervention can be applied. Specifically, the implicit capture of potential confounders can be achieved by integrating the causal intervention mechanism into the attention module through cross-sample attention or self-annotation [35, 16, 32, 36]. Additionally, Liu et al. proposed event-level causal visual question answering utilizing visual front-door intervention, which uses attention to aggregate local and global causality-aware visual-linguistic representations [15]. Therefore, causal inference has achieved remarkable performance in cross-modal tasks [11]. Compared with the previous works that address VQA or image captioning, we focus on a more challenging task MRG, which requires modeling complex visual-linguistic interaction in medical images and textual reports [37, 30, 38]. And we propose a visual-linguistic front-door causal intervention that simultaneously eliminates spurious correlations from visual and linguistic confounders.
II-C Medical Report Generation
@@
Recently, MRG methods have followed the works of image captioning and have shown remarkable performance. However, the abnormal descriptions and lesion regions in patient samples only constitute a small portion, leading to the visual-linguistic bias in the MRG task. To address this issue, knowledge-aware methods prevail because they can explore and distill the knowledge to accurately identify abnormalities and describe them using appropriate terminology. Specifically, the bias from language can be mitigated by utilizing the general graph [6] as prior knowledge [7, 39] However, fixed and limited knowledge is insufficient to address complex medical problems. Therefore, Yang et al. [9] and Li et al. [40] employed RadGraph [41] for knowledge retrieval and dynamic knowledge construction. Additionally, utilizing a well-trained label extractor or manually annotated region labels for classification tasks, and further incorporating them to assist in report generation, can also be considered as approaches leveraging external knowledge [42, 43, 3].
Although the knowledge-based model is a promising direction, acquiring knowledge is expensive and challenging to transfer to other tasks. Therefore, CA [8] and CMCL [44] explored the potential abnormal regions by comparing them with normal samples. Moreover, knowledge can be retrieved or constructed using templates corresponding to similar images [9, 40], which effectively mitigates the issue of representation distortion caused by excessive focus on knowledge. Additionally, the template-based approach requires several samples for comparison, memory-driven transformers can implicitly store the visual [45] and linguistic [5] context without templates. However, considering only the context would lead to cross-modal bias. Thus, the models align the cross-modal information in the embedding space and incorporate it as memory, which can be retrieved by visual-linguistic features [46, 43]. The visual-linguistic bias hinders the robustness and reliability of MRG. To mitigate the cross-modal bias, we introduce a lightweight VLCI approach that implicitly mitigates cross-modal confounders, uncovers the true cross-modal causality through causal front-door intervention, and reduces the need for additional annotation when discovering abnormalities.
III Method
@@
In this section, we first introduce the visual-linguistic pre-training (VLP), followed by the two essential cross-modal causal intervention modules, i.e., the Visual Deconfounding Module (VDM) and the Linguistic Deconfounding Module (LDM). Then, we introduce how to integrate VDM and LDM into the VLCI for cross-modal causal intervention.
@@ III-A Overview
A typical MRG model can generate a series of medical findings and insights given a medical image. As shown in Fig. 2, upon receiving the Chest X-ray image , the VLCI first utilizes a visual feature extractor to obtain , which guides the generation of the next word from the prefix text embedding . However, due to the presence of confounders in both the visual and linguistic modalities, the non-causal model may capture the spurious correlations between “Normal Heart” and “No Effusion”, which causes the neglect of “Pleural Effusion” accompanied by “Normal Heart”, as shown in Fig. 2(a). Fortunately, the SCM of VLCI in Fig. 2(a), which shows the link between and can be cut off by the mediators. The existence of mediators enables the intervention and adjustment of the relation between and , thereby revealing the true causal relation between the two variables. Thus, we perform a causal intervention on the multi-modal feature prior to the decoder to estimate the deconfounded probability of the correct word. This estimation is achieved by aggregating the probabilities of each sub-distribution of mediators, as illustrated in Fig. 2(a). Nevertheless, due to the absence of elaborate datasets and a well-trained feature extractor, we conduct a causal front-door intervention to eliminate the spurious correlations via mediator implicitly. However, the estimation of both confounders and the mediator requires sufficient prior information, i.e., various visual-linguistic concepts. Therefore, we leverage the Visual-Language Pre-training (VLP) model to construct the correlation between the visual contexts and linguistic concepts before conducting the causal intervention.
III-B Visual-Linguistic Pre-training (VLP)

### v5

# Cross-Modal Causal Representation Learning for Radiology Report Generation

@@ Abstract
Radiology Report Generation (RRG) is essential for computer-aided diagnosis and medication guidance, which can relieve the heavy burden of radiologists by automatically generating the corresponding radiology reports according to the given radiology image. However, generating accurate lesion descriptions remains challenging due to spurious correlations from visual-linguistic biases and inherent limitations of radiological imaging, such as low resolution and noise interference. To address these issues, we propose a two-stage framework named Cross-Modal Causal Representation Learning (CMCRL), consisting of the Radiological Cross-modal Alignment and Reconstruction Enhanced (RadCARE) pre-training and the Visual-Linguistic Causal Intervention (VLCI) fine-tuning. In the pre-training stage, RadCARE introduces a degradation-aware masked image restoration strategy tailored for radiological images, which reconstructs high-resolution patches from low-resolution inputs to mitigate noise and detail loss. Combined with a multiway architecture and four adaptive training strategies (e.g., text postfix generation with degraded images and text prefixes), RadCARE establishes robust cross-modal correlations even with incomplete data. In the VLCI phase, we deploy causal front-door intervention through two modules: the Visual Deconfounding Module (VDM) disentangles local-global features without fine-grained annotations, while the Linguistic Deconfounding Module (LDM) eliminates context bias without external terminology databases. Experiments on IU-Xray and MIMIC-CXR show that our CMCRL pipeline significantly outperforms state-of-the-art methods, with ablation studies confirming the necessity of both stages. Code and models are available at https://github.com/WissingChen/CMCRL.
@@
Radiology Report Generation, causality, visual-language pre-training.I Introduction
@@
Radiology images are widely used in clinical procedures [1], providing significant evidence for disease analysis and radiological diagnosis [2]. Nevertheless, observing suspicious lesions and writing a coherent diagnosis report is time-consuming, even for experienced radiologists. Furthermore, inexperienced radiologists often fail to capture tiny abnormalities due to the high requirement for clinical knowledge. To relieve this issue, Radiology Report Generation (RRG) has emerged and attracted growing interest in recent years [3]. RRG extracts features from radiology images and generates the corresponding reports, which is similar to image captioning [4]. However, the current RRG faces three challenges that are significantly different from the image captioning task: 1) Longer sentence generation (60-100 tokens) tends to accumulate a larger bias, whereas image captions typically consist of fewer than 20 tokens [5], 2) The necessity to capture all key regions in the radiology image (i.e., abnormalities and diagnostic evidence) and a low tolerance for factually incorrect information [3], 3) More complex linguistic and visual semantic patterns require a proficient understanding of radiological information, whereas entities in natural images are diverse and easily distinguishable [6, 2]. Therefore, these challenges impose significant limitations on modeling visual-linguistic interactions and learning informative cross-modal representations for accurate RRG [5, 7].
@@
To tackle the aforementioned challenges, current RRG methods have made significant efforts, such as the memory-driven module for longer sentence generation [5], additional knowledge for more accurate description [8], and contrast with normal samples (i.e., images without lesion areas) for the capture of abnormalities (i.e., lesion areas within images) [9]. Most of the previous RRG methods aim to capture the latent subtle differences in images (visual biases) and learn a concise set of key descriptions in the text (linguistic biases) for accurate long-sequence generation. However, these methods usually focus on training computationally expensive models based on a large number of region-level annotations [3] and task-specific knowledge111These methods build the template or knowledge database laboriously, making it hard to transfer those approaches directly to other datasets [10]., rather than focusing on identifying and mitigating biases inherent in the data itself.
@@
As shown in Fig 1 (a-b), the significant visual and linguistic biases in numerous image-text data build a spurious correlation between “Normal Heart” and “No Effusion”, leading to the incorrect and unreliable reports. To mitigate the cross-modal bias and encourage the model to infer based on the corresponding visual basis, as shown in Fig 1 (c), causal inference [11] has shown promising performance in several visual-linguistic tasks [12]. However, directly applying existing causal methods to the RRG task may yield unsatisfactory results due to the unobservable bias in the visual and linguistic domain and the complex visual-linguistic interaction in radiology images and textual reports. Although back-door intervention can cut off the shortcut path, it requires approximating the observable bias using a well-trained visual object detector or a well-constructed linguistic dictionary. Fortunately, causal front-door intervention gives a feasible way to mitigate visual-linguistic spurious correlations without the calculation of unobservable bias. With causal front-door intervention, we can eliminate the spurious cross-modal correlations effectively by introducing an additional mediator [13, 14] and generate an accurate description of “Pleural Effusion”. The mediator can be assumed as the feature of “Pleural” in different findings (e.g., “Normal Heart”) to estimate the sub-distribution of “Pleural Effusion” [15]. However, the accurate and reasonable acquisition of the mediator is challenging, especially without the support of additional radiological knowledge.
To effectively obtain reliable estimates of mediator, well-aligned radiological multi-modal representations are crucial. However, existing multi-modal approaches are limited by singular pre-training strategies and fail to fully leverage data with missing modalities within datasets [16]. Moreover, given the issues of low-resolution inputs and noise interference present in radiological data, directly adopting multi-modal pre-training models designed for natural image tasks does not effectively capture and align multi-modal features [17, 18]. This further restricts the performance of subsequent causal interventions.
@@
Motivated by the characteristic multi-modal radiological data and effectiveness of causal inference in deconfounding the cross-modal bias, we introduce a Cross-Modal Causal Representation Learning (CMCRL) framework for RRG. This framework includes a pre-training model specifically designed for radiological data, termed Radiological Cross-modal Alignment and Reconstruction Enhanced (RadCARE), and a lightweight cross-modal causal intervention model without requiring an observable bias assumption, named Visual-Linguistic Causal Intervention (VLCI), to mitigate biases in visual and linguistic data. Our RadCARE framework introduces a degradation-aware masked image restoration strategy that reconstructs high-resolution patches from degraded inputs, explicitly mitigating detail loss while preserving anatomical coherence. This is coupled with a multi-way architecture enabling four adaptive training strategies, including postfix text generation via prefix text or multi-modal data, and masked image restoration via degraded images or combined with complete text. Our VLCI framework contains the visual deconfounding module (VDM) and linguistic deconfounding module (LDM) based on the causal front-door intervention paradigm. In VDM, the visual mediator is constructed by local detail information (e.g., lung texture) and global contour (e.g., lung contour) from radiology images, to disentangle the visual features. The linguistic confounders can be eliminated by the LDM, which estimates the change in the probability of word embedding caused by visual details and linguistic context. In summary, our main contributions are listed as follows:
@@
To alleviate the problem of unpaired data and capture detailed features when pre-training cross-modal data, we propose RadCARE that integrates degradation-aware masked image restoration with postfix text generation for pre-training in various data situations (e.g., unpaired, single modality), which is efficient and easy to implement.
To mitigate cross-modal biases, we propose visual-linguistic causal intervention modules VDM and LDM, integrated into the VLCI without requiring additional data or a well-trained semantic extractor for guidance.
@@
We propose a CMCRL framework for RRG, which introduces mediators without additional knowledge, to deconfound the visual-linguistic features by causal front-door intervention. To the best of our knowledge, we are the first to conduct a cross-modal causal intervention for RRG. Experiments show that we achieve state-of-the-art performance on IU-Xray and MIMIC-CXR datasets.
@@ II-A Image Captioning
Image captioning, based on the encoder-decoder framework, aims to describe image content in text. Recent advancements have enhanced performance in three areas: visual representation learning, linguistic generation, and training strategies [19]. While CNN-based global feature extraction compresses information excessively [20], integrating visual saliency through regional features and combining visual-semantic data improves results [21]. In linguistic processing, coarse-to-fine captioning [22] and multi-layer LSTM integration [23] have shown promise but are limited by training efficiency. Transformer-based models [24] leverage self-attention for modality learning and cross-attention for multi-modal integration, significantly boosting performance [25, 26]. Performance is further improved by training strategies such as sample order optimization [27], reinforcement learning [4], and visual-linguistic pre-training [28, 18].
With the proliferation of Large Language Models (LLMs), image captioning achieved significant breakthroughs in both model architecture and generation capabilities. The V2L-Tokenizer [29] model demonstrated SOTA performance across multiple tasks through a non-finetuning approach. Meanwhile, prompt learning offered new insights into the controllability of image captioning [30, 31]. Additionally, models like GPT-4o[32] and Qwen2-VL[33] integrated advanced visual processing capabilities with powerful language generation, achieving remarkable results without extensive fine-tuning. Compared to image captioning, RRG shares similar structures [19] but focuses on subtle lesion detection in radiology images, generating longer and more complex descriptions.
@@ II-B Causal Inference
Causality offers a robust approach to mitigating the visual-linguistic bias stemming from the heterogeneity of multi-modal data by eliminating spurious correlations [11, 12]. Causal inference [12] estimates hidden causal effects within distributions, significantly improving model generalization by addressing confounders through back-door, front-door, or counterfactual interventions. This method enhances performance in tasks like image classification [34], semantic segmentation [35], visual feature representation [36, 37], image captioning [26], and visual question answering (VQA)[13]. For instance, Wang et al.[38] applied back-door intervention to improve Faster R-CNN, which subsequently enhanced performance in VQA [39] and image captioning [26]. Additionally, Liu et al. proposed an event-level causal VQA model using front-door intervention, which leverages attention to integrate local and global causality-aware visual-linguistic representations [13]. This approach also addresses confounding by simulating causal interventions based on human priors [40].
Since confounders are often unobservable, front-door and counterfactual interventions are particularly effective. These interventions can capture hidden confounders by integrating causal mechanisms into attention modules through cross-sample attention or self-annotation [41, 37, 42]. In the era of LLMs, causal learning has emerged as a key direction for enhancing both the generalization ability and interpretability. On one hand, by integrating causality into the processes of pre-training, fine-tuning, and inference, LLMs can better handle complex tasks, while also improving their performance in multi-modal scenarios [43, 44]. On the other hand, causal learning enables LLMs to move beyond mere statistical correlations and understand the underlying causal mechanisms in the data, thereby generating more reliable and ethically aligned outputs [45]. Unlike prior work focusing on VQA or image captioning, we tackle the more complex task of RRG, which involves modeling intricate visual-linguistic interactions in radiology images and textual reports [46, 35, 47]. Our proposed front-door causal intervention method addresses and eliminates both visual and linguistic spurious correlations.
II-C Radiology Report Generation
@@
Recently, RRG methods, inspired by image captioning, have achieved notable success. However, abnormal descriptions and lesion regions make up only a small portion of patient samples, leading to visual-linguistic bias. Knowledge-aware approaches address this by leveraging external knowledge to identify abnormalities and describe them accurately [6, 8]. Although fixed knowledge is limited for complex radiological cases, dynamic knowledge construction using RadGraph [48] has been proposed [10, 49]. Furthermore, external knowledge can be integrated through label extractors or manually annotated regions to enhance report generation [50, 51, 3].
While knowledge-based models show promise, acquiring knowledge is costly and difficult to generalize. To mitigate this, CA [9], CMCL [52], CAMANet [53] explored abnormal regions by comparing them with normal samples. Templates based on similar images also help reduce representation distortion [10, 49]. Memory-driven transformers can store visual [54, 55] and linguistic [5] contexts without templates but risk cross-modal bias. To address this, models align cross-modal data in embedding spaces for retrieval [56, 51]. With the advancement of multi-modal large language models (MLLMs), their robust visual reasoning capabilities, internal knowledge, and language generation skills have significantly enhanced performance in radiology report generation tasks. However, the inherent tendency of large models to produce hallucinations continues to pose challenges for this task. Consequently, some methods have focused on developing approaches that enable large models to generate descriptions consistent with the findings in radiological images [57, 58, 59]. To further reduce cross-modal bias, we introduce a lightweight VLCI approach, which applies causal front-door intervention to uncover true cross-modal causality and reduce reliance on additional annotations.
III Methodology
@@
In this section, we first introduce the RadCARE framework for pre-training, followed by the two essential cross-modal causal intervention modules, i.e., the Visual Deconfounding Module (VDM) and the Linguistic Deconfounding Module (LDM). Then, we introduce how to integrate VDM and LDM into the VLCI for cross-modal causal intervention and complete the CMCRL framework.
@@ III-A Overview
A typical RRG model can generate a series of radiological findings and insights given a radiological image. Upon receiving the Chest X-ray image , the model first utilizes a visual feature extractor to obtain , which guides the generation of the next word from the prefix text embedding , as shown in Fig. 2 (a). However, due to the presence of confounders in both the visual and linguistic modalities, the non-causal model may capture the spurious correlations between “Normal Heart” and “No Effusion”, which causes the neglect of “Pleural Effusion” accompanied by “Normal Heart”, as shown in Fig. 2 (c). Fortunately, the mediators can cut off the link between and , and is denoted the multi-modal feature. Mediators enable the intervention and adjustment of the relation between and , thereby revealing the true causal relation between the two variables. Thus, we perform a causal intervention on the multi-modal feature prior to the decoder to estimate the deconfounded probability of the correct word, as shown in Fig. 2 (b). Nevertheless, due to the absence of elaborate datasets and a well-trained feature extractor, we conduct a causal front-door intervention to eliminate the spurious correlations via mediators . However, the estimation of both confounders and the mediator requires sufficient prior information, i.e., various visual-linguistic concepts. Therefore, we leverage the RadCARE to construct the correlation between the visual contexts and linguistic concepts before conducting the causal intervention using VLCI.
III-B RadCARE